1. A dog walks 12 meters to the north and then 16 meters back to the south.

For this motion, what is the distance moved? _________________

What is the magnitude and direction of the displacement? _________________

2. A lady shopper walks eastward 3.2 meters and then westward 7.2 meters.

For this motion, what is the distance moved? _________________

What is the magnitude and direction of the displacement? _________________

3. A basketball player runs 2.5 meters northward and then back pedals 6.8 meters southward.

For this motion, what is the distance moved? _________________

What is the magnitude and direction of the displacement? _________________


1
Expert's answer
2021-11-14T17:17:53-0500

1. A dog walks 12 meters to the north and then 16 meters back to the south.

For this motion, what is the distance moved? 12+16=28 m

What is the magnitude and direction of the displacement? 16-12=4 m South

2. A lady shopper walks eastward 3.2 meters and then westward 7.2 meters.

For this motion, what is the distance moved? 3.2+7.2=10.4 m.

What is the magnitude and direction of the displacement? 7.2-3.2=4 m West

3. A basketball player runs 2.5 meters northward and then back pedals 6.8 meters southward.

For this motion, what is the distance moved? 2.5+6.8=9.3 m

What is the magnitude and direction of the displacement? 6.8-2.5=4.3 m.